# Reviving Thorup’s Shortcut Conjecture

### v1

@@
Aaron Bernstein New York University    Henry Fleischmann Carnegie Mellon University    George Z. Li22footnotemark: 2    Bernhard Haeupler INSAIT, Sofia University “St. Kliment Ohridski” and ETH Zürich    Maximilian Probst Gutenberg ETH Zürich    Gary Hoppenworth University of Michigan    Seth Pettie55footnotemark: 5    Thatchaphol Saranurak55footnotemark: 5    Yonggang Jiang MPI-INF and Saarland University    Leon Schiller Hasso Plattner Institute, University of Potsdam
@@ 1.2 Contributions
For example, if $V_{+}=\emptyset$ (there are no Steiner vertices) then the thickness is 11, and all prior lower bounds [Hes03, HP21, LVWX22, VXX24, HXX25] are merely for 1-thick shortcut graphs. A 2-thick shortcut graph is one for which the in- and out-neighbors of $V_{+}$-vertices are all original $V$-vertices.
@@ 2.1 Warm-up Construction — Graph Family $\mathcal{G}_{1}$
We construct a graph family $\mathcal{G}_{1}$ that does not quite imply polynomial diameter lower bounds for $\tilde{O}(m)$-size shortcut sets. This will motivate some of the design choices introduced in the graph family $\mathcal{G}_{k}$ in Section $2.$
@@
Vertex set $V$ and paths $\mathcal{P}$ have a very natural geometric interpretation as an arrangement of points and lines in ${\mathbb{R}}^{d+1}$. Notice that the vertices of each path $P_{s,t}\in\mathcal{P}$ lie on a line of the form $\{x\cdot(1,\vec{v})+(0,p)\mid x\in{\mathbb{R}}\}$ in ${\mathbb{R}}^{d+1}$ for $\vec{v}\in\mathcal{B}(r,d)$ and $(0,p)\in S$. Moreover, because $\mathcal{B}(r,d)$ consists only of extreme points, each such direction $\vec{v}$ cannot be expressed as a convex combination of other vectors in $\mathcal{B}(r,d)$. This geometric perspective is useful for interpreting the Uniqueness and Disjointness properties of Lemma $2.$
@@ 2.2 Hesse’s $\mathcal{G}_{k}$ Family
Section 2.2 has a similar geometric interpretation as Lemma $2.$ The main difference is that the division into subdirections allows for two critical paths to have small overlap, but they still cannot overlap in two layers at distance $k$, which implies the third property. See Section D.1 for formal proof.
@@
  • The number of critical pairs/critical paths is $|P|=\Theta\left(\frac{N\Delta^{k}r^{dk}}{(r+1)^{dk}}\right)$ by Section $2.$
@@ 3 Necessity of Thick Steiner Shortcuts
We now introduce a noisy variant of the graph family $\mathcal{G}_{k}$, which we call $\widetilde{\cal{G}}_{k}(c,d,p)$. It is defined as $\widetilde{\cal{G}}_{k}(c,d,p)=\{\widetilde{G}(p)\mid G\in\mathcal{G}_{k}(c,d)\}$, where $\widetilde{G}(p)$ denotes the (random) graph obtained from a given $G$ when deleting each edge independently with probability pp. $\widetilde{\cal{G}}_{k}(c,d,p)$ is thus a family of random graphs that we obtain from the elements of $\mathcal{G}_{k}(c,d)$ by deleting each edge independently with probability pp. Here $p$ should not be thought of as an absolute constant, but a function of $c,d,r$ that is inversely proportional to the diameter of $G$.
@@ 3.1 Basic Properties
Let $G\in\mathcal{G}_{k}$ and $G_{+}$ be a $t$-thick Steiner shortcut for $G$ with $M$ edges and diameter $h$. There is a $t$-thick Steiner shortcut $G^{\prime}=(V\cup V^{\prime},E\cup E^{\prime})$ for $G$ such that $E^{\prime}$ is the union of less than $(\ell k)^{2}$ normalized $t$-thick shortcut sets, each with at most $M$ edges. The diameter of $G^{\prime}$ is less than hkhk.
@@
In this subsection, we improve the bounds from Section 3.2 under the assumption that edges in $G_{k}$ are deleted independently with probability $p=1/(k\ell)$. That is, we consider the (random) graphs from the family $\widetilde{\cal{G}}(c,d,p)$, where in this section $c=3,d=20000$. The introduction of this type of “noise” implies that each critical path remains present with constant probability but at the same time it limits the maximum efficiency of Steiner shortcuts which enables improved bounds on their thickness. Precisely, we show that after noising, no sufficiently efficient Steiner shortcut can survive. To this end, given any $A\subseteq L_{i},B\subseteq L_{j}$, we relate the number of critical paths going from AA to $B$ to the number of edges on these paths (i.e. their support size). Concretely, defining $\sigma:=\max\{|A|,|B|\}$, we show that whenever there are $\gg\sigma^{3/2}\mathrm{polylog}(n)$ critical paths going from AA to $B$, their support size is so large that with high probability at least one edge supporting our paths is deleted when applying noise. Using a union bound, this in turn implies that the number of critical paths between any pair of sets $A\subseteq V_{i},B\subseteq V_{j}$ is at most $\sigma^{3/2}$, which is much smaller than the trivial bound of $\sigma^{2}$. This limits the maximum efficiency any Steiner shortcut of a given thickness for $G\in\widetilde{\cal{G}}_{k}(c,d,p)$ with high probability, and allows us to rule out shortcuts of thickness $(8/7)k$ instead of just $k$ as in Theorem 3.1.
Formally, our proof relies on the following two lemmas, the first of which connects the number of critical paths between AA and $B$ to their support size in a $G\in\mathcal{G}_{k}(c,d)$.
@@
Consider any $G\in\mathcal{G}_{k}(c,d)$ and let $i,j$ be such that $|i-j|\geq 4rk$. Futher, let $A\subseteq L_{i},B\subseteq L_{j}$ be two subsets of vertices in layers $i$ and $j$, respectively. Then, if there is a set $\mathcal{P}$ of at least $\frac{s^{3/2}}{2\sqrt{rk}}$ critical paths that connect a vertex in AA to a vertex in $B$, the number of edges supported by the paths in $\mathcal{P}$ between the layers $L_{i}$ and $L_{j}$ is at least $s$.
@@
Assume that the number of critical paths from AA to $B$ is at least $\eta s$ for some suitable $\eta$ ($\eta$ can be seen as the efficiency of a star shortcutting AA to $B$). Suppose that the support size of these paths at most $s$ across the layers $L_{i},L_{i+1},\ldots,L_{j}$. We show that a contradiction arises if $\eta\geq\frac{\sqrt{s}}{2\sqrt{rk}}$. This shows that many critical paths between $A,B$ imply a large support size.
Denote the set of critical paths with one vertex in AA and one in $B$ by $\mathcal{P}$. Note that our assumption on the support size of the paths in $\mathcal{P}$ implies that the average number of outgoing edges associated to any path in $\mathcal{P}$ per layer in between AA and $B$ in $\widetilde{G}$ is at most s/|ij|s/|i-j|, so there must be a layer $L_{t},t\in[i,j-1]$ with at most $s/(4rk)$ supported outgoing edges incident to $L_{t}$. We let $U_{1}$ be the set of vertices in $L_{t}$ that are on a path in $\mathcal{P}$ and note that $|U_{1}|\leq s/(4rk)$ because the number of supported vertices is at most the number of supported outgoing edges of $L_{t}$.
@@
Note that the vertices corresponding to $N_{H}(u)$ have a pairwise Euclidean distance within the dkdk–dimensional grid associated to layer $U$ of at most $2rk$ since all of them are reachable from $u$ by following a critical direction, each of which has total Euclidean norm $\leq rk$. We claim that following all critical paths incident to the vertices in $N_{H}(u)$ for >2rk>2rk steps, we reach $p(u)$ distinct vertices. Indeed, if there were a vertex $v$ reachable from two distinct critical paths starting at two vertices $u_{1},u_{2}\in N_{H}(u)$ after >2rk>2rk steps, then $\|\vec{u_{1}}-\vec{u_{2}}\|_{2}>2rk$ contradicting the fact that the Euclidean distance between $u_{1},u_{2}$ is at most $2rk$ as required by the assumption that $u_{1},u_{2}$ are both neighbors of $u$ in HH. [^6] For any $u\in L_{i}$, let $\vec{u}$ denote the position of $u$ in the dkdk-dimensional integer grid associated with layer $L_{i}$. This uses that the critical directions are distinct vectors with integer coordinates. Accordingly, if $|i-j|\geq 4rk$, then
@@
as otherwise, we obtain a contradiction to the assumption that all paths are supported by at most $s$ edges. Rearranging the above yields that that whenever $\eta>\frac{\sqrt{s}}{2\sqrt{rk}}$, there is some layer on which all critical paths connecting AA to $B$ are supported by at least $s$ edges. ∎
With the above, $(k\ell\sigma\log(n))^{\frac{3}{2}}$ critical paths between AA and $B$ yield a support size of at least $k\ell\sigma\log(n)$. Using that each edge is deleted independently with probability $p=1/(k\ell)$ then implies that, for every such set, at least one critical path fails.
@@
Consider any $G\in G_{k}$ and let $i,j$ be such that $|i-j|\geq 4rk$. Let $A\subseteq L_{i},B\subseteq L_{j}$ and define $\sigma\coloneqq\max\{|A|,|B|\}$. Then with probability $1-1/\mathrm{poly}(n)$, for all such $i,j,A,B$ in $\widetilde{G}$ with at least $(100\sigma\log(n)\ell k)^{3/2}$ critical paths connecting a vertex in AA to a vertex in $B$, there is at least one such critical path that fails when passing from $G$ to the noised version $\widetilde{G}$.
@@
Let $\gamma:=100\log n$. We show that for every choice of $A\subseteq L_{i},B\subseteq L_{j}$ with more than $(\gamma\sigma k\ell)^{3/2}$ critical paths, at least one path will fail after applying noise. Invoking Lemma 3.9, the support size of these paths between layers $L_{i}$ and $L_{j}$ is at least $\gamma\sigma k\ell$. Furthermore, the number of all $A,B$ with $\max\{|A|,|B|\}=\sigma$ is at most $n^{2\sigma}$, and the probability that for fixed sets $A,B$ no critical path connecting AA to $B$ fails is at most
@@
because we have chosen $p=\frac{1}{k\ell}$. Applying a union bound yields that the probability that there is some $A\subseteq L_{i},B\subseteq L_{j}$ with $|i-j|\geq 4rk$ and at least $(\gamma\sigma k\ell)^{3/2}$ critical paths connecting AA and $B$ is at most
@@
This concludes that with probability $1-1/\mathrm{poly}(n)$, there is no pair of sets $A\subseteq L_{i},B\subseteq L_{j}$ such that $|i-j|\geq 4rk$ and such that there are more than $(100s\log(n)\ell k)^{3/2}$ critical paths connecting AA and $B$.
Proof of Theorem $3.$
@@
On the other hand, we know that the critical $E^{*}$-respecting shortest paths passing through $v$ correspond to critical paths passing through some $A\subseteq L_{i}$ and some $B\subseteq L_{j}$ where $L_{i},L_{j}$ are the layers shortcut by the component $C$ of $E^{*}$ that $v$ belongs to. Due to the maximum degree restriction in $E^{*}$, and the fact that $v\in M$, we know that $\sigma:=\max\{|A|,|B|\}\leq\Delta^{(\frac{t}{2}+1)(1+\epsilon^{\prime})}$. Applying Lemma 3.10, we get that the number of critical paths with one vertex in AA and the other in $B$ is $O((k\ell\sigma\log(n))^{3/2})=O(\Delta^{\frac{3t}{4}(1+\epsilon^{\prime})+\epsilon^{\prime}+6\epsilon+1})=O(\Delta^{\frac{3t}{4}(1+2\epsilon^{\prime})})$ since $\ell=O(r^{3})$ and $r=O(\Delta^{2\epsilon})$. In total, we get that the following inequality must be true
@@ 4.1 Formalization
To formalize this, we need notation for multi-commodity flows. Let $G=(V,E)$ be a graph with edge capacity $c:E\rightarrow\mathbb{R}_{\geq 0}$ and edge length $\ell:E\rightarrow\mathbb{R}_{\geq 0}$. A (multi-commodity) flow $F$ in $G$ assigns value $F(P)\geq 0$ for each directed path $P$ in $G$. We say that $P$ is a flow path of $F$ if $F(P)>0$. The congestion on edge $e$ of $F$ is $\mathrm{cong}_{F}(e)=\frac{\sum_{e\in P}F(P)}{c(e)}$ and the congestion of $F$ is $\mathrm{cong}_{F}=\max_{e\in E}\mathrm{cong}_{F}(e)$. The length of $F$ is the maximum total length overall flow paths of $F$, i.e., $\max_{P:F(P)>0}\ell(P)$. The hopbound of $F$ is the maximum number of edges overall flow paths of $F$, i.e., $\max_{P:F(P)>0}|E(P)|$. For any vertex set $U$, a demand $D$ on $U$ assigns a value $D(u,v)\geq 0$ for each vertex pair $(u,v)\in U$. The demand $D_{F}$ routed by $F$ is defined as $D_{F}(u,v):=\sum_{(u,v)\text{-path }P}F(P)$. We say that $D$ is routable with length $\ell$, congestion $\kappa$, and hopbound $h$ if there exists a flow $F$ with length $\ell$, congestion $\kappa$, and hopbound $h$ that routes $D$. When $D$ is routable in $G$ with congestion 11, we sometimes just say that $D$ is routable in $G$. With this, we are ready to define LC-flow shortcuts.
@@
    Every demand on $V$ routable in $G$ with length $\lambda$ and congestion 11 is routable in $G^{\prime}$ with length $\lambda s$, congestion 11, and hopbound $h$, and
@@
    Every demand on $V$ routable in $G^{\prime}$ with length $\lambda$ and congestion 11 is routable in $G$ with length $\lambda$ and congestion $\kappa$.

### v2

@@
Aaron Bernstein New York University    Henry Fleischmann Carnegie Mellon University    Maximilian Probst Gutenberg ETH Zürich    Bernhard Haeupler INSAIT, Sofia University “St. Kliment Ohridski” and ETH Zürich    Gary Hoppenworth University of Michigan    Yonggang Jiang MPI-INF and Saarland University    George Z. Li22footnotemark: 2    Seth Pettie55footnotemark: 5    Thatchaphol Saranurak55footnotemark: 5    Leon Schiller Hasso Plattner Institute, University of Potsdam
@@ 1.2 Contributions
For example, if $V_{+}=\emptyset$ (there are no Steiner vertices) then the thickness is $1$, and all prior lower bounds [Hes03, HP21, LVWX22, VXX24, HXX25] are merely for 1-thick shortcut graphs. A 2-thick shortcut graph is one for which the in- and out-neighbors of $V_{+}$-vertices are all original $V$-vertices.
@@ 2.1 Warm-up Construction — Graph Family $\mathcal{G}_{1}$
We construct a graph family $\mathcal{G}_{1}$ that does not quite imply polynomial diameter lower bounds for $\tilde{O}(m)$-size shortcut sets. This will motivate some of the design choices introduced in the graph family $\mathcal{G}_{k}$ in Section 2.2.
@@
Vertex set $V$ and paths $\mathcal{P}$ have a very natural geometric interpretation as an arrangement of points and lines in ${\mathbb{R}}^{d+1}$. Notice that the vertices of each path $P_{s,t}\in\mathcal{P}$ lie on a line of the form $\{x\cdot(1,\vec{v})+(0,p)\mid x\in{\mathbb{R}}\}$ in ${\mathbb{R}}^{d+1}$ for $\vec{v}\in\mathcal{B}(r,d)$ and $(0,p)\in S$. Moreover, because $\mathcal{B}(r,d)$ consists only of extreme points, each such direction $\vec{v}$ cannot be expressed as a convex combination of other vectors in $\mathcal{B}(r,d)$. This geometric perspective is useful for interpreting the Uniqueness and Disjointness properties of Lemma 2.2.
@@ 2.2 Hesse’s $\mathcal{G}_{k}$ Family
Section 2.2 has a similar geometric interpretation as Lemma 2.2. The main difference is that the division into subdirections allows for two critical paths to have small overlap, but they still cannot overlap in two layers at distance $k$, which implies the third property. See Section D.1 for formal proof.
@@
  • The number of critical pairs/critical paths is $|P|=\Theta\left(\frac{N\Delta^{k}r^{dk}}{(r+1)^{dk}}\right)$ by Section 2.2.
@@ 3 Necessity of Thick Steiner Shortcuts
We now introduce a noisy variant of the graph family $\mathcal{G}_{k}$, which we call $\widetilde{\cal{G}}_{k}(c,d,p)$. It is defined as $\widetilde{\cal{G}}_{k}(c,d,p)=\{\widetilde{G}(p)\mid G\in\mathcal{G}_{k}(c,d)\}$, where $\widetilde{G}(p)$ denotes the (random) graph obtained from a given $G$ when deleting each edge independently with probability $p$. $\widetilde{\cal{G}}_{k}(c,d,p)$ is thus a family of random graphs that we obtain from the elements of $\mathcal{G}_{k}(c,d)$ by deleting each edge independently with probability $p$. Here $p$ should not be thought of as an absolute constant, but a function of $c,d,r$ that is inversely proportional to the diameter of $G$.
@@ 3.1 Basic Properties
Let $G\in\mathcal{G}_{k}$ and $G_{+}$ be a $t$-thick Steiner shortcut for $G$ with $M$ edges and diameter $h$. There is a $t$-thick Steiner shortcut $G^{\prime}=(V\cup V^{\prime},E\cup E^{\prime})$ for $G$ such that $E^{\prime}$ is the union of less than $(\ell k)^{2}$ normalized $t$-thick shortcut sets, each with at most $M$ edges. The diameter of $G^{\prime}$ is less than $hk$.
@@
In this subsection, we improve the bounds from Section 3.2 under the assumption that edges in $G_{k}$ are deleted independently with probability $p=1/(k\ell)$. That is, we consider the (random) graphs from the family $\widetilde{\cal{G}}(c,d,p)$, where in this section $c=3,d=20000$. The introduction of this type of “noise” implies that each critical path remains present with constant probability but at the same time it limits the maximum efficiency of Steiner shortcuts which enables improved bounds on their thickness. Precisely, we show that after noising, no sufficiently efficient Steiner shortcut can survive. To this end, given any $A\subseteq L_{i},B\subseteq L_{j}$, we relate the number of critical paths going from $A$ to $B$ to the number of edges on these paths (i.e. their support size). Concretely, defining $\sigma:=\max\{|A|,|B|\}$, we show that whenever there are $\gg\sigma^{3/2}\mathrm{polylog}(n)$ critical paths going from $A$ to $B$, their support size is so large that with high probability at least one edge supporting our paths is deleted when applying noise. Using a union bound, this in turn implies that the number of critical paths between any pair of sets $A\subseteq V_{i},B\subseteq V_{j}$ is at most $\sigma^{3/2}$, which is much smaller than the trivial bound of $\sigma^{2}$. This limits the maximum efficiency any Steiner shortcut of a given thickness for $G\in\widetilde{\cal{G}}_{k}(c,d,p)$ with high probability, and allows us to rule out shortcuts of thickness $(8/7)k$ instead of just $k$ as in Theorem 3.1.
Formally, our proof relies on the following two lemmas, the first of which connects the number of critical paths between $A$ and $B$ to their support size in a $G\in\mathcal{G}_{k}(c,d)$.
@@
Consider any $G\in\mathcal{G}_{k}(c,d)$ and let $i,j$ be such that $|i-j|\geq 4rk$. Futher, let $A\subseteq L_{i},B\subseteq L_{j}$ be two subsets of vertices in layers $i$ and $j$, respectively. Then, if there is a set $\mathcal{P}$ of at least $\frac{s^{3/2}}{2\sqrt{rk}}$ critical paths that connect a vertex in $A$ to a vertex in $B$, the number of edges supported by the paths in $\mathcal{P}$ between the layers $L_{i}$ and $L_{j}$ is at least $s$.
@@
Assume that the number of critical paths from $A$ to $B$ is at least $\eta s$ for some suitable $\eta$ ($\eta$ can be seen as the efficiency of a star shortcutting $A$ to $B$). Suppose that the support size of these paths at most $s$ across the layers $L_{i},L_{i+1},\ldots,L_{j}$. We show that a contradiction arises if $\eta\geq\frac{\sqrt{s}}{2\sqrt{rk}}$. This shows that many critical paths between $A,B$ imply a large support size.
Denote the set of critical paths with one vertex in $A$ and one in $B$ by $\mathcal{P}$. Note that our assumption on the support size of the paths in $\mathcal{P}$ implies that the average number of outgoing edges associated to any path in $\mathcal{P}$ per layer in between $A$ and $B$ in $\widetilde{G}$ is at most s/|ij|s/|i-j|, so there must be a layer $L_{t},t\in[i,j-1]$ with at most $s/(4rk)$ supported outgoing edges incident to $L_{t}$. We let $U_{1}$ be the set of vertices in $L_{t}$ that are on a path in $\mathcal{P}$ and note that $|U_{1}|\leq s/(4rk)$ because the number of supported vertices is at most the number of supported outgoing edges of $L_{t}$.
@@
Note that the vertices corresponding to $N_{H}(u)$ have a pairwise Euclidean distance within the dkdk–dimensional grid associated to layer $U$ of at most $2rk$ since all of them are reachable from $u$ by following a critical direction, each of which has total Euclidean norm $\leq rk$. We claim that following all critical paths incident to the vertices in $N_{H}(u)$ for >2rk>2rk steps, we reach $p(u)$ distinct vertices. Indeed, if there were a vertex $v$ reachable from two distinct critical paths starting at two vertices $u_{1},u_{2}\in N_{H}(u)$ after >2rk>2rk steps, then $\|\vec{u_{1}}-\vec{u_{2}}\|_{2}>2rk$ contradicting the fact that the Euclidean distance between $u_{1},u_{2}$ is at most $2rk$ as required by the assumption that $u_{1},u_{2}$ are both neighbors of $u$ in HH. [^6] For any $u\in L_{i}$, let $\vec{u}$ denote the position of $u$ in the $dk$-dimensional integer grid associated with layer $L_{i}$. This uses that the critical directions are distinct vectors with integer coordinates. Accordingly, if $|i-j|\geq 4rk$, then
@@
as otherwise, we obtain a contradiction to the assumption that all paths are supported by at most $s$ edges. Rearranging the above yields that that whenever $\eta>\frac{\sqrt{s}}{2\sqrt{rk}}$, there is some layer on which all critical paths connecting $A$ to $B$ are supported by at least $s$ edges. ∎
With the above, $(k\ell\sigma\log(n))^{\frac{3}{2}}$ critical paths between $A$ and $B$ yield a support size of at least $k\ell\sigma\log(n)$. Using that each edge is deleted independently with probability $p=1/(k\ell)$ then implies that, for every such set, at least one critical path fails.
@@
Consider any $G\in G_{k}$ and let $i,j$ be such that $|i-j|\geq 4rk$. Let $A\subseteq L_{i},B\subseteq L_{j}$ and define $\sigma\coloneqq\max\{|A|,|B|\}$. Then with probability $1-1/\mathrm{poly}(n)$, for all such $i,j,A,B$ in $\widetilde{G}$ with at least $(100\sigma\log(n)\ell k)^{3/2}$ critical paths connecting a vertex in $A$ to a vertex in $B$, there is at least one such critical path that fails when passing from $G$ to the noised version $\widetilde{G}$.
@@
Let $\gamma:=100\log n$. We show that for every choice of $A\subseteq L_{i},B\subseteq L_{j}$ with more than $(\gamma\sigma k\ell)^{3/2}$ critical paths, at least one path will fail after applying noise. Invoking Lemma 3.9, the support size of these paths between layers $L_{i}$ and $L_{j}$ is at least $\gamma\sigma k\ell$. Furthermore, the number of all $A,B$ with $\max\{|A|,|B|\}=\sigma$ is at most $n^{2\sigma}$, and the probability that for fixed sets $A,B$ no critical path connecting $A$ to $B$ fails is at most
@@
because we have chosen $p=\frac{1}{k\ell}$. Applying a union bound yields that the probability that there is some $A\subseteq L_{i},B\subseteq L_{j}$ with $|i-j|\geq 4rk$ and at least $(\gamma\sigma k\ell)^{3/2}$ critical paths connecting $A$ and $B$ is at most
@@
This concludes that with probability $1-1/\mathrm{poly}(n)$, there is no pair of sets $A\subseteq L_{i},B\subseteq L_{j}$ such that $|i-j|\geq 4rk$ and such that there are more than $(100s\log(n)\ell k)^{3/2}$ critical paths connecting $A$ and $B$.
Proof of Theorem 3.3.
@@
On the other hand, we know that the critical $E^{*}$-respecting shortest paths passing through $v$ correspond to critical paths passing through some $A\subseteq L_{i}$ and some $B\subseteq L_{j}$ where $L_{i},L_{j}$ are the layers shortcut by the component $C$ of $E^{*}$ that $v$ belongs to. Due to the maximum degree restriction in $E^{*}$, and the fact that $v\in M$, we know that $\sigma:=\max\{|A|,|B|\}\leq\Delta^{(\frac{t}{2}+1)(1+\epsilon^{\prime})}$. Applying Lemma 3.10, we get that the number of critical paths with one vertex in $A$ and the other in $B$ is $O((k\ell\sigma\log(n))^{3/2})=O(\Delta^{\frac{3t}{4}(1+\epsilon^{\prime})+\epsilon^{\prime}+6\epsilon+1})=O(\Delta^{\frac{3t}{4}(1+2\epsilon^{\prime})})$ since $\ell=O(r^{3})$ and $r=O(\Delta^{2\epsilon})$. In total, we get that the following inequality must be true
@@ 4.1 Formalization
To formalize this, we need notation for multi-commodity flows. Let $G=(V,E)$ be a graph with edge capacity $c:E\rightarrow\mathbb{R}_{\geq 0}$ and edge length $\ell:E\rightarrow\mathbb{R}_{\geq 0}$. A (multi-commodity) flow $F$ in $G$ assigns value $F(P)\geq 0$ for each directed path $P$ in $G$. We say that $P$ is a flow path of $F$ if $F(P)>0$. The congestion on edge $e$ of $F$ is $\mathrm{cong}_{F}(e)=\frac{\sum_{e\in P}F(P)}{c(e)}$ and the congestion of $F$ is $\mathrm{cong}_{F}=\max_{e\in E}\mathrm{cong}_{F}(e)$. The length of $F$ is the maximum total length overall flow paths of $F$, i.e., $\max_{P:F(P)>0}\ell(P)$. The hopbound of $F$ is the maximum number of edges overall flow paths of $F$, i.e., $\max_{P:F(P)>0}|E(P)|$. For any vertex set $U$, a demand $D$ on $U$ assigns a value $D(u,v)\geq 0$ for each vertex pair $(u,v)\in U$. The demand $D_{F}$ routed by $F$ is defined as $D_{F}(u,v):=\sum_{(u,v)\text{-path }P}F(P)$. We say that $D$ is routable with length $\ell$, congestion $\kappa$, and hopbound $h$ if there exists a flow $F$ with length $\ell$, congestion $\kappa$, and hopbound $h$ that routes $D$. When $D$ is routable in $G$ with congestion $1$, we sometimes just say that $D$ is routable in $G$. With this, we are ready to define LC-flow shortcuts.
@@
    Every demand on $V$ routable in $G$ with length $\lambda$ and congestion $1$ is routable in $G^{\prime}$ with length $\lambda s$, congestion $1$, and hopbound $h$, and
@@
    Every demand on $V$ routable in $G^{\prime}$ with length $\lambda$ and congestion $1$ is routable in $G$ with length $\lambda$ and congestion $\kappa$.